# Nonsmooth Calabi–Yau structures for algebras and coalgebras

### v1

@@ 1. Introduction
between augmented dg-$k$-algebras and conilpotent dg-$k$-coalgebras, for a fixed field $k$. The functors in question are given by the bar and cobar constructions. This equivalence in particular underlies the modern approach to noncommutative derived deformation theory [49, §3]. DG coalgebras can also be thought of as pseudocompact dg algebras (or more accurately, their continuous linear duals) and we switch between these perspectives when convenient.
@@
We remark that one can use this theorem to give a Morita duality result, stating that if $A$ is a smooth local Calabi–Yau dg algebra then ${{\mathbf{pvd}}}\kern 1.0pt(A)$ is a right Calabi–Yau dg category (see 5.35 and 8.16 for the details).
It will be key for us to work with nonsmooth dg algebras (and dually, non-proper dg coalgebras). For example, the dual numbers $k[\varepsilon]/\varepsilon^{2}$ is a nonsmooth Calabi–Yau dg algebra, which leads to the surprising conclusion that its Koszul dual, the power series ring $k\llbracket x\rrbracket$, is a pseudocompact symmetric algebra. One can also prove this directly from the definitions (3.12); there is no contradiction here as the ‘pseudocompact linear dual’ is not computed as the linear dual of the underlying vector space. More generally, we prove that the pseudocompact algebra of power series $k\llbracket x_{1},\ldots,x_{n}\rrbracket$, with $x_{i}$ in degree $d_{i}$, is $(n-\sum_{i}d_{i})$-symmetric (8.7). In a loose sense this can be thought of as a derived local complete intersection property for the associated formal derived scheme.
@@
Our main application is a new characterisation of Poincaré duality spaces. For a fixed field $k$ and pointed topological space $X$, we let $C_{\bullet}X$ be the dg coalgebra of $k$-chains on $X$ and $C^{\bullet}X$ the dg algebra of $k$-cochains. When $Y=\Omega X$ is a loop space, recall that concatenation of loops makes $C_{\bullet}Y$ into a dg algebra (in fact a dg Hopf algebra).
Theorem B (9.30 and 9.40).
@@
Our notion of Poincaré duality space is a fully derived version that gives chain-level Poincaré duality for all local systems of dg-$k$-vector spaces. In particular, it is sensitive to the base field: we warn that a space $X$ could be a Poincaré duality space for a given choice of $k$ and not for a different choice. Theorem B generalises a characterisation of simply connected Poincaré duality spaces due to Felix, Halperin, and Thomas [18]; namely that they are Gorenstein spaces. In fact we show that a Poincaré duality space is always a Gorenstein space, and that the converse is true for simply connected finite CW complexes (9.42). In order to prove Theorem B, we make use of the dg Hopf algebra structure on $C_{\bullet}\Omega X$. This use of the Hopf structure is key in order to translate between the one-sided conditions (Gorenstein, Frobenius) and the two-sided conditions (Calabi–Yau, symmetric). Specifically, we show that certain dg Hopf algebras are Calabi–Yau if and only they are Gorenstein (9.8); the proof rests on a computation of their Hochschild cohomology, which generalises a theorem of Ginzburg and Kumar [24]. We also give an application to unimodular Lie algebras (9.14). As applications of our results on Poincaré duality, we also obtain some results in string topology, rational homotopy theory, and Lie group (co)homology.
@@
While preparing this paper, we learned of the similar results obtained by Holstein and Rivera [35]. They obtain analogous Koszul duality statements for left and right Calabi–Yau dg categories. There are three key differences between [35] and the current paper: first is that we drop any smoothness/properness hypotheses completely. This means that we are unable to obtain any results about (negative) cyclic homology, since in our setting CY structures are no longer given by Hochschild classes. However it means that we are able to get wider classes of results: for example, we show that the pseudocompact algebra $k\llbracket x\rrbracket$ is Frobenius (= weak right CY), a notion which is only defined in [35] for proper pseudocompact algebras. For example, our Theorem A generalises the weak one-object cases of [35, 3.26] and [35, 3.33].
@@
The authors would like to thank Callum Galvin, Timothy De Deyn, Isambard Goodbody, John Greenlees, Julian Holstein, Sebastian Opper, David Pauksztello, Alex Takeda, and Michael Wemyss for helpful conversations.
@@ 2.2. Comodules and contramodules
If $C$ is a coalgebra, its linear dual $C^{*}$ is an algebra under the convolution product $f.g=m(f\otimes g)\Delta$, where $m:k\otimes k\to k$ is the multiplication isomorphism. In fact, $C^{*}$ is a pseudocompact algebra, and the linear dual is a contravariant equivalence between the category of coalgebras and the category of pseudocompact algebras. Similarly, the linear dual gives a contravariant equivalence between the categories of $C$-comodules and pseudocompact $C^{*}$-modules. One can equivalently phrase all of our theorems about coalgebras in the language of pseudocompact algebras.
@@
The existence of $\square_{C}$ is [55, 1.2.4] and associativity is [55, 1.2.5]. The existence of $\odot_{C}$ is [55, 5.1.1, 5.1.2] and the cotensor-contratensor associativity is [55, 5.2, Proposition 1]. The existence of $\mathrm{Hom}_{C}$ is [55, 5.1.2].
For brevity we will write $(U,V)\coloneqq\mathrm{Hom}_{k}(U,V)$. Let $T$ be a $C$-comodule and $Z$ a $C$-contramodule. Recall that $\mathrm{Cohom}_{C}(T,Z)$ is defined as the coequaliser of the natural diagram $(C\otimes T,Z)\rightrightarrows(T,Z)$ given by the $C$-coaction on $T$ and the $C$-contraaction on $Z$ [55, 3.2.1]. Suppose that $T$ is a $D$-$C$-bicomodule and $Z$ is an $E$-$C$-bicontramodule. We wish to construct a map $(D\otimes E,\mathrm{Cohom}_{C}(T,Z))\to\mathrm{Cohom}_{C}(T,Z)$ exhibiting $\mathrm{Cohom}_{C}(T,Z)$ as an $E$-$D$-bicontramodule. Since $\mathrm{Hom}(D\otimes E,-)$ commutes with coequalisers, it suffices to show that the coequaliser diagram defining $\mathrm{Cohom}_{C}(T,Z)$ is a diagram of $D$-$E$-bicontramodules. Let $\rho:T\to D\otimes T$ and $\sigma:(E,Z)\to Z$ denote the structure maps. Then the $D$-$E$-contraaction on $(T,Z)$ is given by the composition
@@
and similarly for $(C\otimes T,Z)$. Moreover the two maps in the coequaliser are $D$-$E$-bicontralinear, since they only use the $C$-co/contraactions. Hence $\mathrm{Cohom}_{C}(T,Z)$ is a $D$-$E$-bicontramodule, as required. The cohom-cotensor adjunction is [56, 5.2, Proposition 2].
@@
Since $(D,-)$ preserves equalisers, $\mathrm{Hom}^{C}(W,Y)$ is hence a $D$-contramodule. The cohom-contratensor adjunction is [56, 5.2, Proposition 3]. ∎
@@ 3.1. One-sided duality
The existence of the model structures and Quillen functors is [56, §8]. Everything else is contained in [56, §6.5]. ∎
The Quillen functors between ${\mathbf{Mod}}\text{-}A$ and ${\mathbf{Comod}}\text{-}C$ are given by twisted tensor products, and the functors between ${\mathbf{Mod}}\text{-}A$ and ${\mathbf{Ctrmod}}\text{-}C$ are given by twisted Homs [56, §6.2]. Specifically, let $\tau:C\to A$ denote the associated twisting cochain. If $M$ is an $A$-module then the corresponding $C$-comodule is the twisted tensor product $M\widetilde{\otimes}C$; the underlying graded vector space is $M\otimes C$ and the differential is twisted by the term $m\otimes c\mapsto m\tau(c_{1})\otimes c_{2}$, where we use Sweedler notation for the comultiplication of $C$. The coaction on $M\widetilde{\otimes}C$ is given by the natural coaction of $C$ on itself. Similarly, if $N$ is a $C$-comodule then the corresponding $A$-module is $N\widetilde{\otimes}A$, where the differential is twisted by the term $n\otimes a\mapsto n_{1}\otimes\tau(n_{2})a$, where we again use Sweedler notation for the coaction $N\to N\otimes C$. Note that $M\widetilde{\otimes}C$ is a cofree comodule.
@@ 3.3. Linear duals for comodules
We refer to $M^{*}$ as the linear dual of $M$. We warn that $M^{*}$ depends on the $C$-comodule structure of $M$, and not just on the underlying vector space, since limits in $C$-comodules are not computed by the underlying vector spaces:
@@
If $M$ is a $C^{e}$-comodule we denote the $C^{e}$-comodule $M^{*}$ by $D(M)$, to avoid confusion between one- and two-sided linear duals; as before the underlying $C^{\circ}$-comodule of $D(M)$ need not be $M^{*}$.
@@ 3.4. One- and two-sided duals for contramodules
Say that a $C$-comodule $M$ is homotopy quasi-finite (hqf for short) if, for all finite dimensional comodules $N$, the vector space ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{C}(N,M)$ is an element of ${{\mathbf{pvd}}}\kern 1.0pt(k)$. This is a homotopical version of Takeuchi’s definition of a quasi-finite comodule [62]. Let ${{\mathbf{hqf}}^{\mathrm{co}}}(C)$ denote the full subcategory of $D^{\mathrm{co}}(C)$ on the hqf comodules. Note that we have an equivalence ${{\mathbf{hqf}}^{\mathrm{co}}}(C)\simeq{{\mathbf{pvd}}}\kern 1.0pt\left(D^{\mathrm{co}}(C)\right)$. If $(A,C)$ is a Koszul duality pair, then across Koszul duality the category ${{\mathbf{hqf}}^{\mathrm{co}}}(C)$ corresponds to ${{\mathbf{pvd}}}\kern 1.0pt(A)$: this is because finite dimensional comodules correspond to perfect $A$-modules, and $V\in{{\mathbf{pvd}}}\kern 1.0pt(A)$ if and only if ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{A}({{\mathbf{Perf}}}\kern 1.0pt(A),V)\in{{\mathbf{pvd}}}\kern 1.0pt(k)$. Similarly, say that a contramodule $X$ is hqf precisely when the corresponding comodule $X\odot_{C}C$ is; the hqf contramodules form a full subcategory ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C)$ of $D^{\mathrm{ctr}}(C)$ that can be identified with ${{\mathbf{pvd}}}\kern 1.0ptD^{\mathrm{ctr}}(C)$. Since the linear dual is an equivalence ${{\mathbf{pvd}}}\kern 1.0pt(A)^{\circ}\to{{\mathbf{pvd}}}\kern 1.0pt(A^{\circ})$, by 3.17 it follows that the functor ${\mathrm{\mathbb{R}}}\mathrm{Hom}^{C}(-,C^{*})$ gives a contravariant equivalence between ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C)^{\circ}$ and ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C^{\circ})$.
@@
Letting $A^{\prime}$ be the algebra $\Omega(C^{e})$, the Koszul duality pair $(C^{e},A^{\prime})$ gives by 3.17 a commutative diagram
@@ 4.2. Twists
The following proposition will allow us to deduce the existence of Nakayama automorphisms for derived Frobenius algebras (for the discrete version, see [47, §16E]).
@@ 5.1. Properness and locality for algebras
If (1) holds then by [58, 5.7] combined with 5.4 we see that (2) holds. Since (2) clearly implies (1) we see that (1) and (2) are equivalent. Moreover, certainly (3) implies (1) so we just need to show that (1) implies (3). To do this, certainly if (1) holds then $A$ is proper, so we just need to prove that it is local. By 5.4 we only need to prove that the finite dimensional algebra $H^{0}A$ is local. But by assumption it is an iterated extension between copies of $k$, and hence must be local. ∎
@@ 5.2. Smoothness and regularity for algebras
Claim $(1)$ is standard: if $M\in{{\mathbf{pvd}}}\kern 1.0ptA$ and $P$ is a perfect $A$-bimodule then then $M\otimes^{\mathbb{L}}_{A}P\in{{\mathbf{Perf}}}\kern 1.0ptA$; it is enough to show this for $P=A^{e}$ which is clear. In particular, if $A$ is smooth then we may take $P=A$. Claim (2) follows from a theorem of Orlov [54, 1.11]. ∎
@@
If $A$ is smooth then it is regular. Conversely, if $A$ is regular then assume without loss of generality that $A$ is finite dimensional. Then it is smooth by [54, 3.12]. ∎

### v2

@@ 1. Introduction
between augmented dg-$k$-algebras and conilpotent dg-$k$-coalgebras, for a fixed field $k$. The functors in question are given by the bar and cobar constructions. This equivalence in particular underlies the modern approach to noncommutative derived deformation theory [51, §3]. DG coalgebras can also be thought of as pseudocompact dg algebras (or more accurately, their continuous linear duals) and we switch between these perspectives when convenient.
@@
We remark that one can use this theorem to give a Morita duality result, stating that if $A$ is a smooth local Calabi–Yau dg algebra then ${{\mathbf{pvd}}}\kern 1.0pt(A)$ is a right Calabi–Yau dg category (see 5.35 and 8.17 for the details).
It will be key for us to work with nonsmooth dg algebras (and dually, non-proper dg coalgebras). For example, the dual numbers $k[\varepsilon]/\varepsilon^{2}$ is a nonsmooth Calabi–Yau dg algebra, which leads to the surprising conclusion that its Koszul dual, the power series ring $k\llbracket x\rrbracket$, is a pseudocompact symmetric algebra. One can also prove this directly from the definitions (3.12); there is no contradiction here as the ‘pseudocompact linear dual’ is not computed as the linear dual of the underlying vector space. More generally, we prove that the pseudocompact graded algebra of power series $k\llbracket x_{1},\ldots,x_{n}\rrbracket$, with $x_{i}$ in degree $d_{i}$, is $(n-\sum_{i}d_{i})$-symmetric (8.7). In a loose sense this can be thought of as a derived local complete intersection property for the associated formal derived scheme. In fact, a similar argument shows that a discrete commutative complete local Gorenstein $k$-algebra also has this self-duality property (8.10), which is in some sense a manifestation of Matlis duality for zero-dimensional Gorenstein rings.
@@
Our main application is a new characterisation of Poincaré duality spaces. For a fixed field $k$ and pointed topological space $X$, we let $C_{\bullet}X$ be the dg coalgebra of $k$-chains on $X$ and $C^{\bullet}X$ the dg algebra of $k$-cochains. When $Y=\Omega X$ is a loop space, recall that concatenation of loops makes $C_{\bullet}Y$ into a dg algebra (in fact - up to quasi-isomorphism - a dg Hopf algebra).
Theorem B (9.32 and 9.42).
@@
Our notion of Poincaré duality space is a fully derived version that gives chain-level Poincaré duality for all local systems of dg-$k$-vector spaces. In particular, it is sensitive to the base field: we warn that a space $X$ could be a Poincaré duality space for a given choice of $k$ and not for a different choice. Theorem B generalises a characterisation of simply connected Poincaré duality spaces due to Felix, Halperin, and Thomas [18]; namely that they are Gorenstein spaces. In fact we show that a Poincaré duality space is always a Gorenstein space, and that the converse is true for simply connected finite CW complexes (9.44). In order to prove Theorem B, we make use of the dg Hopf algebra structure on $C_{\bullet}\Omega X$ (strictly, one needs to take normalised chains on the Kan loop group of $\mathrm{Sing}X$). This use of the Hopf structure is key in order to translate between the one-sided conditions (Gorenstein, Frobenius) and the two-sided conditions (Calabi–Yau, symmetric). Specifically, we show that certain dg Hopf algebras are Calabi–Yau if and only they are Gorenstein (9.8); the proof rests on a computation of their Hochschild cohomology, which generalises a theorem of Ginzburg and Kumar [24]. We also give an application to unimodular Lie algebras (9.15). As applications of our results on Poincaré duality, we also obtain some results in string topology, rational homotopy theory, and Lie group (co)homology.
@@
While preparing this paper, we learned of the similar results obtained by Holstein and Rivera [37]. They obtain analogous Koszul duality statements for left and right Calabi–Yau dg categories. There are three key differences between [37] and the current paper: first is that we drop any smoothness/properness hypotheses completely. This means that we are unable to obtain any results about (negative) cyclic homology, since in our setting CY structures are no longer given by Hochschild classes. However it means that we are able to get wider classes of results: for example, we show that the pseudocompact algebra $k\llbracket x\rrbracket$ is Frobenius (= weak right CY), a notion which is only defined in [37] for proper pseudocompact algebras. For example, our Theorem A generalises the weak one-object cases of [37, 3.26] and [37, 3.33].
@@
The authors would like to thank Benjamin Briggs, Timothy De Deyn, Callum Galvin, Isambard Goodbody, John Greenlees, Julian Holstein, Sebastian Opper, David Pauksztello, Manuel Rivera, Alex Takeda, and Michael Wemyss for helpful conversations.
@@ 2.2. Comodules and contramodules
If $C$ is a coalgebra, its linear dual $C^{*}$ is an algebra under the convolution product $f.g=m(f\otimes g)\Delta$, where $m:k\otimes k\to k$ is the multiplication isomorphism. In fact, $C^{*}$ is a pseudocompact algebra, and the linear dual is a contravariant equivalence between the category of coalgebras and the category of pseudocompact algebras. Similarly, the linear dual gives a contravariant equivalence between the categories of $C$-comodules and pseudocompact $C^{*}$-modules. One can equivalently phrase all of our theorems about coalgebras in the language of pseudocompact algebras. In this paper we will often pass between the coalgebraic language and the pseudocompact language without comment.
@@
The existence of $\square_{C}$ is [57, 1.2.4] and associativity is [57, 1.2.5]. The existence of $\odot_{C}$ is [57, 5.1.1, 5.1.2] and the cotensor-contratensor associativity is [57, 5.2, Proposition 1]. The existence of $\mathrm{Hom}_{C}$ is [57, 5.1.2].
For brevity we will write $(U,V)\coloneqq\mathrm{Hom}_{k}(U,V)$. Let $T$ be a $C$-comodule and $Z$ a $C$-contramodule. Recall that $\mathrm{Cohom}_{C}(T,Z)$ is defined as the coequaliser of the natural diagram $(C\otimes T,Z)\rightrightarrows(T,Z)$ given by the $C$-coaction on $T$ and the $C$-contraaction on $Z$ [57, 3.2.1]. Suppose that $T$ is a $D$-$C$-bicomodule and $Z$ is an $E$-$C$-bicontramodule. We wish to construct a map $(D\otimes E,\mathrm{Cohom}_{C}(T,Z))\to\mathrm{Cohom}_{C}(T,Z)$ exhibiting $\mathrm{Cohom}_{C}(T,Z)$ as an $E$-$D$-bicontramodule. Since $\mathrm{Hom}(D\otimes E,-)$ commutes with coequalisers, it suffices to show that the coequaliser diagram defining $\mathrm{Cohom}_{C}(T,Z)$ is a diagram of $D$-$E$-bicontramodules. Let $\rho:T\to D\otimes T$ and $\sigma:(E,Z)\to Z$ denote the structure maps. Then the $D$-$E$-contraaction on $(T,Z)$ is given by the composition
@@
and similarly for $(C\otimes T,Z)$. Moreover the two maps in the coequaliser are $D$-$E$-bicontralinear, since they only use the $C$-co/contraactions. Hence $\mathrm{Cohom}_{C}(T,Z)$ is a $D$-$E$-bicontramodule, as required. The cohom-cotensor adjunction is [58, 5.2, Proposition 2].
@@
Since $(D,-)$ preserves equalisers, $\mathrm{Hom}^{C}(W,Y)$ is hence a $D$-contramodule. The cohom-contratensor adjunction is [58, 5.2, Proposition 3]. ∎
@@ 3.1. One-sided duality
The existence of the model structures and Quillen functors is [58, §8]. Everything else is contained in [58, §6.5]. ∎
The Quillen functors between ${\mathbf{Mod}}\text{-}A$ and ${\mathbf{Comod}}\text{-}C$ are given by twisted tensor products, and the functors between ${\mathbf{Mod}}\text{-}A$ and ${\mathbf{Ctrmod}}\text{-}C$ are given by twisted Homs [58, §6.2]. Specifically, let $\tau:C\to A$ denote the associated twisting cochain. If $M$ is an $A$-module then the corresponding $C$-comodule is the twisted tensor product $M\widetilde{\otimes}C$; the underlying graded vector space is $M\otimes C$ and the differential is twisted by the term $m\otimes c\mapsto m\tau(c_{1})\otimes c_{2}$, where we use Sweedler notation for the comultiplication of $C$. The coaction on $M\widetilde{\otimes}C$ is given by the natural coaction of $C$ on itself. Similarly, if $N$ is a $C$-comodule then the corresponding $A$-module is $N\widetilde{\otimes}A$, where the differential is twisted by the term $n\otimes a\mapsto n_{1}\otimes\tau(n_{2})a$, where we again use Sweedler notation for the coaction $N\to N\otimes C$. Note that $M\widetilde{\otimes}C$ is a cofree comodule.
@@ 3.3. Linear duals for comodules
We refer to $M^{*}$ as the linear dual of $M$. We warn that $M^{*}$ depends on the $C$-comodule structure of $M$, and not just on the underlying vector space, since limits in $C$-comodules are not computed by the underlying vector spaces. For a concrete example of this, see 3.12 below.
[…]
@@
Remark 3.13.
The ‘naïve’ linear dual functor is well defined on the full subcategory ${{\mathbf{fd}}}(C)\hookrightarrow D^{\mathrm{co}}(C)$ of compact objects, and the full linear dual functor of 3.11 can be obtained as a homotopy Kan extension of the naïve linear dual along the inclusion.
If $M$ is a $C^{e}$-comodule we denote the $C^{e}$-comodule $M^{*}$ by $D(M)$, to avoid confusion between one- and two-sided linear duals; as before the underlying $C^{\circ}$-comodule of $D(M)$ need not be $M^{*}$. Once again, we extend this definition to pseudocompact bimodules in the obvious manner.
@@ 3.4. One- and two-sided duals for contramodules
Say that a $C$-comodule $M$ is homotopy quasi-finite (hqf for short) if, for all finite dimensional comodules $N$, the vector space ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{C}(N,M)$ is an element of ${{\mathbf{pvd}}}\kern 1.0pt(k)$. This is a homotopical version of Takeuchi’s definition of a quasi-finite comodule [64]. Let ${{\mathbf{hqf}}^{\mathrm{co}}}(C)$ denote the full subcategory of $D^{\mathrm{co}}(C)$ on the hqf comodules. Note that we have an equivalence ${{\mathbf{hqf}}^{\mathrm{co}}}(C)\simeq{{\mathbf{pvd}}}\kern 1.0pt\left(D^{\mathrm{co}}(C)\right)$. If $(A,C)$ is a Koszul duality pair, then across Koszul duality the category ${{\mathbf{hqf}}^{\mathrm{co}}}(C)$ corresponds to ${{\mathbf{pvd}}}\kern 1.0pt(A)$: this is because finite dimensional comodules correspond to perfect $A$-modules, and $V\in{{\mathbf{pvd}}}\kern 1.0pt(A)$ if and only if ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{A}({{\mathbf{Perf}}}\kern 1.0pt(A),V)\in{{\mathbf{pvd}}}\kern 1.0pt(k)$. Similarly, say that a contramodule $X$ is hqf precisely when the corresponding comodule $X\odot_{C}C$ is; the hqf contramodules form a full subcategory ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C)$ of $D^{\mathrm{ctr}}(C)$ that can be identified with ${{\mathbf{pvd}}}\kern 1.0ptD^{\mathrm{ctr}}(C)$. Since the linear dual is an equivalence ${{\mathbf{pvd}}}\kern 1.0pt(A)^{\circ}\to{{\mathbf{pvd}}}\kern 1.0pt(A^{\circ})$, by 3.18 it follows that the functor ${\mathrm{\mathbb{R}}}\mathrm{Hom}^{C}(-,C^{*})$ gives a contravariant equivalence between ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C)^{\circ}$ and ${{\mathbf{hqf}}^{\mathrm{ctr}}}(C^{\circ})$.
@@
Letting $A^{\prime}$ be the algebra $\Omega(C^{e})$, the Koszul duality pair $(C^{e},A^{\prime})$ gives by 3.18 a commutative diagram
@@ 4.2. Twists
The following proposition will allow us to deduce the existence of Nakayama automorphisms for derived Frobenius algebras (for the discrete version, see [49, §16E]).
@@ 5.1. Properness and locality for algebras
If (1) holds then by [60, 5.7] combined with 5.4 we see that (2) holds. Since (2) clearly implies (1) we see that (1) and (2) are equivalent. Moreover, certainly (3) implies (1) so we just need to show that (1) implies (3). To do this, certainly if (1) holds then $A$ is proper, so we just need to prove that it is local. By 5.4 we only need to prove that the finite dimensional algebra $H^{0}A$ is local. But by assumption it is an iterated extension between copies of $k$, and hence must be local. ∎
@@ 5.2. Smoothness and regularity for algebras
Claim $(1)$ is standard: if $M\in{{\mathbf{pvd}}}\kern 1.0ptA$ and $P$ is a perfect $A$-bimodule then then $M\otimes^{\mathbb{L}}_{A}P\in{{\mathbf{Perf}}}\kern 1.0ptA$; it is enough to show this for $P=A^{e}$ which is clear. In particular, if $A$ is smooth then we may take $P=A$. Claim (2) follows from a theorem of Orlov [56, 1.11]. ∎
@@
If $A$ is smooth then it is regular. Conversely, if $A$ is regular then assume without loss of generality that $A$ is finite dimensional. Then it is smooth by [56, 3.12]. ∎